# Traffic-R1: Reinforced LLMs Bring Human-Like Reasoning to Traffic Signal Control Systems

### v1

@@ Abstract.
Traffic signal control (TSC) is vital for mitigating congestion and sustaining urban mobility. In this paper, we introduce Traffic-R1, a foundation model with human-like reasoning for TSC systems. Our model is developed through self-exploration and iteration of reinforced large language models (LLMs) with expert guidance in a simulated traffic environment. Compared to traditional reinforcement learning (RL) and recent LLM-based methods, Traffic-R1 offers three significant advantages. First, Traffic-R1 delivers zero-shot generalisation, transferring unchanged to new road networks and out-of-distribution incidents by utilizing its internal traffic control policies and human-like reasoning. Second, its 3B-parameter architecture is lightweight enough for real-time inference on mobile-class chips, enabling large-scale edge deployment. Third, Traffic-R1 provides an explainable TSC process and facilitates multi-intersection communication through its self-iteration and a new synchronous communication network. Extensive benchmarks demonstrate that Traffic-R1 sets a new state of the art, outperforming strong baselines and training-intensive RL controllers. In practice, the model now manages signals for more than 55,000 drivers daily, shortening average queues by over 5% and halving operator workload. Our checkpoint is available at https://huggingface.co/Season998/Traffic-R1.
Traffic signal control, large language model, reinforcement learning
copyright: none
1. Introduction
Figure 1. Introduction of Traffic-R1, a foundation (covering six features) reinforced LLM for TSC systems.
@@ 1. Introduction
Rapid urbanization and surging vehicle ownership intensify congestion, wasting billions of productive hours, burning vast fuel reserves, and driving nearly a quarter of urban greenhouse emissions. Prolonged delays raise crash rates, slow emergency response, exacerbate cardiopulmonary pollution, and unfairly burden transit-poor communities. Within this broad societal context, traffic-signal control (TSC), which coordinates phase sequences and durations at signalized intersections, remains a principal lever for mitigating congestion and improving network throughput (Yau et al., 2017; Ye et al., 2019; Wei et al., 2021).
@@
Traditional controllers such as FixedTime (Koonce et al., 2008) and MaxPressure (Varaiya, 2013) rely on fixed heuristics and thus adapt poorly to fluctuating demand. Reinforcement learning (RL) replaces these hand-crafted rules with a data-driven policy: each cycle observes lane queues, delays, and neighboring signal states, selects a phase (or duration) as the action, and receives a reward linked to delay reduction or throughput gain. Deep RL further augments this paradigm by learning the policy end-to-end with expressive function approximators. These advances achieve impressive benchmark scores in simulation (Srinivasan et al., 2006; Zhao et al., 2011). However, field deployment remains rare as existing methods still suffer from (i) cross-region generalization: policies tuned on one city transfer poorly to another (Zhao et al., 2024; Wei et al., 2021); (ii) interpretability: nontransparent decisions undermine practitioner trust (Vouros, 2022; Glanois et al., 2024); and (iii) robustness to out-of-distribution (OOD) events: models falter during incidents or emergency-vehicle priority scenarios (Glanois et al., 2024; Wei et al., 2018).
@@
Recently, large language models (LLMs) have been enlisted to alleviate these shortcomings. Two integration paradigms dominate. In the LLM Explainer paradigm (Lai et al., 2023; Yuan et al., 2025), an LLM is trained to verbalize the policy of an RL controller, translating opaque action choices into natural language rationales. In contrast, the LLM assistant framework (Wang et al., 2025a, 2024) keeps the RL agent in charge of routine control and consults an LLM only when OOD incidents arise. Both paradigms are promising yet remain distant from large-scale deployment: LLM Explainers inherit the coverage and performance ceiling of the underlying RL policy, and their post-hoc narratives can diverge from the controller’s true internal logic (Wu et al., 2025; Glanois et al., 2024; Wei et al., 2018; Agarwal et al., 2024; Malin et al., 2025); LLM Assistants introduce additional prompt engineering and repeated LLM queries, inflating latency and computation while providing limited benefit for everyday signal timing. Hence practitioners remain cautious about adopting current LLM-enhanced TSC solutions.
@@
To date, operational TSC systems still rely on heuristic rule sets and substantial human oversight to cope with routine flow and unexpected incidents (Mandhare et al., 2018; Müller et al., 2021; Mahavar and Juremalani, 2018). Bridging this research-deployment gap requires a foundational model for TSC systems, i.e., a single, versatile agent capable of (1) zero-shot generalization to unseen traffic networks and OOD incidents, (2) resource-efficient inference on edge hardware such as mobile platforms, and (3) human-like & transparent reasoning that supports explainable decision-making and multi-intersection coordination. Figure 1) schematically illustrates these three desiderata in the envisioned foundational agent.
@@
We answer this call with Traffic-R1, a lightweight reinforced LLM with human-like reasoning capabilities, designed as a foundational TSC model that incorporates all six key features shown in Figure 1 for real-world deployment. Built on Qwen2.5-3B, a efficient LLM optimized for resource-constrained devices, Traffic-R1 employs a two-stage agentic RL finetuning approach to enhance TSC generalization. This includes an offline RL stage, where the model is finetuned using offline TSC recordings and decisions from human experts to integrate their knowledge, and an online RL stage, where it explores dynamic simulated traffic environments to adapt to various scenarios. Drawing on recent studies (Bai et al., 2022; Liu et al., 2024; Yu et al., 2025), Traffic-R1 is trained to generate Chain-of-Thought responses and actions through self-iteration during RL finetuning, guided by a policy-based reward model that includes both format and action rewards. In the offline RL stage, action rewards are calculated based on the differences between the model’s actions and human expert decisions, while in the online RL stage, they are derived from the simulated traffic systems. This approach allows the lightweight Traffic-R1 to develop reasoning and decision-making abilities through self-exploration within human expert decisions and dynamic simulated traffic environments, supporting SOTA zero-shot TSC performance while ensuring resource-efficient inference and robust generalization to OOD scenarios.
Moreover, since all samples for updating parameters are generated by Traffic-R1 itself within our RL finetuning framework, it reduces the risk of degrading general language abilities or experiencing catastrophic forgetting, which can occur due to capacity mismatches between synthetic samples produced by other LLMs and the capabilities of a compact base model. As a result, Traffic-R1 retains strong general language skills alongside explainable human-like reasoning capabilities for transparent decision-making. To further leverage these language abilities, we introduce an asynchronous communication network for LLM-based TSC systems. This network enables LLM agents to communicate and coordinate in a manner similar to human traffic agents, using an asynchronous message-passing mechanism that facilitates effective and transparent coordination across multiple intersections.
In summary, our contributions lie in the following aspects:
@@
  • Foundation model for TSC systems: We present the first LLM-based, general-purpose controller that can operate at any interaction without additional training, handling routine signalling, incident management, and emergency-vehicle prioritization with human-level reasoning.
  • Lightweight yet high-performing: We develop a two-stage agentic RL fine-tuning framework to train a 3B-parameter LLM that outperforms large counterparts (e.g., GPT-4o) and strong RL baselines, while remaining deployable on mobile or edge devices.
@@
  • Human-like reasoning and communication: We achieve human-like reasoning for explainable TSC through self-iteration of the reinforced LLM with policy-based rewards. This reasoning, combined with language capabilities, is further utilized in our asynchronous communication network for effective coordination across multiple intersections.
@@
  • Extensive validation and field deployment: We evaluate Traffic-R1 on standard TSC benchmarks and OOD incident tasks, where it attains stable, state-of-the-art (SOTA) performance. In live deployment on a platform that manages signals for more than 55,000 drivers daily, parallel trials show a 5% cut in average queues and a greater than 50% reduction in operator workload for phase planning and incident response.
2. Related Works
@@
Traffic Signal Control. TSC are essential for effective traffic management. The FixedTime method, one of the earliest and most widely approaches in TSC systems, uses predetermined cycle lengths and phase allocations set by human experts for each intersection (Yau et al., 2017; Serafini and Ukovich, 1989). While this method is straightforward and stable, it is inefficient across dynamic traffic conditions and costs a large amount of human effort (Serafini and Ukovich, 1989; Thunig et al., 2019). The Maxpressure approach introduces adaptive control for better performance by prioritizing vehicle movement based on pressure, defined as the number of queuing vehicles in each direction (Varaiya, 2013; Mercader et al., 2020). However, both fixed-time and Maxpressure methods are limited by their reliance on fixed plans or rules, which struggle to adapt to varying traffic scenarios (Rasheed et al., 2020; Wei et al., 2021; Abdulhai et al., 2003). Advances in machine learning have led to the development of RL-based TSC methods, such as CoLight (Wei et al., 2019), CosLight (Ruan et al., 2024), and MPLight (Chen et al., 2020). These RL approaches have significantly improved TSC performance in simulated environments but often fail to meet the requirements for real-world deployment (Chen et al., 2022; Wei et al., 2021; Qadri et al., 2020). Recent research has explored the integration of LLMs to enhance TSC systems; however, various challenges remain for their real-world deployment. (Lai et al., 2023; Yuan et al., 2025; Wang et al., 2025a, 2024).
RL finetuning for LLM. LLMs often require finetuning to adapt to specific tasks or align with human preferences (Wu et al., 2025; Minaee et al., 2024). In recent years, instruction finetuning (Chung et al., 2024; Dettmers et al., 2023) has gained popularity due to its simplicity and effectiveness in adapting LLMs for various tasks and response formats. In contrast, RL finetuning, particularly Reinforcement Learning from Human Feedback (Ouyang et al., 2022), has been less discussed for downstream applications due to its complex training pipeline and high computational cost (Gao et al., 2023; Bai et al., 2022; Ramamurthy et al., 2022). Recent advances have shown that LLMs can develop reasoning capabilities by themselves through interaction with RL environments. The introduction of Group Policy Reward Optimization (Liu et al., 2024) significantly reduces computational costs by introducing policy-based rewards. RL fine-tuning gradually turns to be popular in serval domains and offers significant advantages (Cao et al., 2024; Xu et al., 2025; Ke et al., 2025): (1) self-exploration reduces the need for extensive training samples, (2) RL guides LLMs to learn abilities rather than imitating or memorize answers, and (3) RL fine-tuning mitigates ability degradation and catastrophic forgetting, as updates to parameters are based on self-generated samples with Kullback-Leibler divergence constraints.
3. Preliminaries
Definition 1.
@@ 3. Preliminaries
Problem Statement 0.
LLM Reasoning for Traffic Signal Control. For each intersection $i$ in the road network, a LLM serves as agent $\pi_{i}$. At time step $t$, agent $\pi_{i}$ receives a textual traffic observation $\mathcal{O}_{i,t}$, incident information $\mathcal{I}_{i,t}$, and messages $\mathcal{M}_{i,t-1}$ from neighboring agents. The TSC problem for intersection $i$ is formulated as optimizing the selection of a signal phase $a_{i,t}\in\mathcal{A}$ and generating reasoning content $\mathcal{R}_{i,t}$ to maximize the reward function, subject to incident-specific and multi-agent coordination constraints. The system dynamics are defined as:
(1) $(\mathcal{R}_{i,t},a_{i,t})=\pi_{i}(\mathcal{O}_{i,t},\mathcal{I}_{i,t},\mathcal{M}_{i,t-1};\theta_{i})$
where $\mathcal{R}_{i,t}$ is the reasoning output of agent $\pi_{i}$ with parameters $\theta_{i}$, $a_{i,t}\in\mathcal{A}$ is the selected phase.
4. Methodology
Figure 2. Introduction of the two-stage RL framework
@@ 4. Methodology
In this section, we describe the training pipeline for Traffic-R1. As shown in Figure 2, we utilize a two-stage RL framework comprising offline human-informed RL and online open-world RL to finetune the LLM. For each stage, we define distinct training datasets, environments, policy optimization methods, and reward designs to support their functionality. Besides, we propose an asynchronous communication network for our reinforced LLM to support efficient multi-intersection coordination in TSC systems.
4.1. Human-Informed Reinforcement Learning
@@
Existing LLM-based methods for TSC typically finetune their models on action trajectories generated by traditional SOTA RL models. While this approach ensures performance in simulated experiments, it has notable limitations. First, finetuning LLMs to imitate traditional RL models confines their performance and generalization to the capabilities of the teacher RL model. Second, action trajectories produced by RL models through iterative optimization in simulated environments may not provide clear reasoning and logic for LLMs, as some RL decision policies and actions can be suboptimal or impractical for general TSC, focusing instead on narrow performance metrics. To address these limitations, we propose a human-informed RL finetuning stage that incorporates an expert-collaborative TSC dataset and offline policy optimization. This approach replaces the RL model teacher with real human traffic experts to guide the finetuning of the LLM.
4.1.1. Expert-Collaborative TSC Dataset Construction
Our RL finetuning approach reduces the training data requirement from hundreds of thousands of samples typically needed for instruction finetuning to just thousands. This enables the creation of a TSC dataset with actions provided by human experts for each traffic scenario. We developed an expert-collaborative TSC dataset consisting of 3,000 high-quality question-answer (QA) samples. As illustrated in Figure 3, we generate diverse traffic observations and their corresponding textual descriptions inspired by (Lai et al., 2023). These are processed by DeepSeek-R1 to produce suggested actions based on the observations. To evaluate the effectiveness of these actions, we utilize the SUMO simulator (Krajzewicz et al., 2012) to model traffic flow variations caused by the suggested actions and assess changes in mean vehicle speed. Additionally, each suggested action is reviewed by two human experts from the traffic control department. For samples where DeepSeek-R1’s suggested actions do not pass validation, human experts manually determine the correct actions. Through this pipeline, we collected 3,000 valid samples with the collaboration of 11 human traffic experts. Notably, we exclude any reasoning content from the dataset and retain only the final actions to encourage the LLM to generate its own reasoning during subsequent RL iterations, rather than imitating external reasoning processes.
Figure 3. Expert-Collaborative Dataset Construction Pipeline
4.1.2. Offline Policy Optimization

### v2

@@ Abstract
We introduce Traffic-R1, a 3B-parameter foundation model with human-like reasoning for Traffic signal control (TSC), developed via self-exploration and iterative reinforcement of LLM with expert guidance in a simulated traffic environment. Compared with traditional reinforcement learning and recent LLM-based methods, Traffic-R1 offers three main advantages: zero-shot generalization, transferring unchanged to new road networks and out-of-distribution incidents by leveraging internal traffic-control policies and reasoning; a compact 3B-parameter design that supports real-time inference on mobile-class chips for edge deployment; and an explainable TSC process that enables multi-intersection coordination through communication and an asynchronous communication network. Extensive benchmarks show Traffic-R1 outperforms strong baselines and training-intensive RL controllers. In production, the model now manages signals affecting over 55,000 drivers daily, reduces average queue lengths by more than 5%, and halves operator workload. Our model is available at https://huggingface.co/Season998/Traffic-R1.
[…]
Traffic-R1: Reinforced LLMs Bring Human-Like Reasoning to Traffic Signal Control Systems
Xingchen Zou1,2, Yuhao Yang3, Zheng Chen2, Xixuan Hao1, Yiqi Chen2, Chao Huang3, Yuxuan Liang1thanks: Corresponding author. Email: yuxliang@outlook.com. 1The Hong Kong University of Science and Technology (Guangzhou), 2PCITECH, 3The University of Hong Kong {xzou428,xhao390}@connect.hkust-gz.edu.cn, {chaohuang75}@gmail.com, {chenzheng1, chenyiqi}@pcitech.com, {yuhao-yang, yuxliang}@outlook.com
1 Introduction
@@
Rapid urbanization and surging vehicle ownership intensify congestion, wasting billions of productive hours, burning vast fuel reserves, and driving nearly a quarter of urban greenhouse emissions. Prolonged delays raise crash rates, slow emergency response, exacerbate cardiopulmonary pollution, and unfairly burden transit-poor communities. Within this broad societal context, traffic-signal control (TSC), which coordinates phase sequences and durations at signalized intersections, remains a principal lever for mitigating congestion and improving network throughput Yau et al. (2017); Ye et al. (2019); Wei et al. (2021); Zhang et al. (2024).
Figure 1: Introduction of Traffic-R1, a foundation (covering six features) reinforced LLM for TSC systems.
@@
Traditional controllers like FixedTime Koonce et al. (2008) and MaxPressure Varaiya (2013) rely on fixed heuristics and thus adapt poorly to fluctuating demand. Reinforcement learning (RL) replaces these hand-crafted rules with a data-driven policy: each cycle observes lane queues, delays, and neighboring signal states, selects a phase (or duration) as the action, and receives a reward linked to delay reduction or throughput gain. Deep RL further augments this paradigm by learning the policy end-to-end with expressive function approximators. These advances achieve impressive benchmark scores in simulation Srinivasan et al. (2006); Zhao et al. (2011). However, field deployment remains rare as existing methods still suffer from (i) cross-region generalization: policies tuned on one city transfer poorly to another Zhao et al. (2024); Wei et al. (2021); (ii) interpretability: nontransparent decisions undermine practitioner trust Vouros (2022); Glanois et al. (2024); and (iii) robustness to out-of-distribution (OOD) events: models falter during incidents or emergency-vehicle priority scenarios Glanois et al. (2024); Wei et al. (2018).
@@
Recently, large language models (LLMs) have been enlisted to alleviate these shortcomings. Two integration paradigms dominate. In the LLM Explainer paradigm Lai et al. (2023); Yuan et al. (2025), an LLM is trained to verbalize the policy of an RL controller, translating opaque action choices into natural language rationales. In contrast, the LLM assistant framework Wang et al. (2025a, 2024) keeps the RL agent in charge of routine control and consults an LLM only when OOD incidents arise. Both paradigms are promising yet remain distant from large-scale deployment: LLM Explainers inherit the coverage and performance ceiling of the underlying RL policy, and their post-hoc narratives can diverge from the controller’s true internal logic Wu et al. (2025); Glanois et al. (2024); Wei et al. (2018); Agarwal et al. (2024); Malin et al. (2025); LLM Assistants introduce additional prompt engineering and repeated LLM queries, inflating latency and computation while providing limited benefit for everyday signal timing. Hence practitioners remain cautious about adopting current LLM-enhanced TSC solutions.
@@
To date, operational TSC systems still rely on heuristic rule sets and substantial human oversight to cope with routine flow and unexpected incidents Mandhare et al. (2018); Müller et al. (2021); Mahavar and Juremalani (2018). Bridging this research-deployment gap requires a foundational model for TSC systems, i.e., a single, versatile agent capable of (1) zero-shot generalization to unseen traffic networks and OOD incidents, (2) resource-efficient inference on edge hardware such as mobile platforms, and (3) human-like & transparent reasoning that supports explainable decision-making and multi-intersection coordination. Figure 1) schematically illustrates these three desiderata in the envisioned foundational agent.
@@
We answer this call with Traffic-R1, a lightweight reinforced LLM with human-like reasoning, designed as a foundational traffic controllers incorporating the six key features from Figure 1. Built on the efficient Qwen2.5-3B, Traffic-R1 employs a two-stage agentic RL finetuning approach: an offline stage integrates human expert knowledge from TSC recordings, and an online stage adapts the model to dynamic scenarios through exploration in simulated environments. Inspired by recent studies Bai et al. (2022); Liu et al. (2024); Yu et al. (2025); Zhou et al. (2023), our RL framework trains Traffic-R1 to generate Chain-of-Thought (CoT) reasoning along with actions via self-iteration. The learning process is guided by a policy-based reward model that incorporates rewards for both output format and action correctness. In the offline stage, action rewards are calculated based on alignment with human expert decisions; in the online stage, they are derived from simulated traffic feedback. This training scheme enables the lightweight model to develop robust reasoning through self-exploration, leading to strong zero-shot TSC performance and generalization to out-of-distribution (OOD) scenarios, all while maintaining resource-efficient inference.
By generating its own training samples within our RL framework, Traffic-R1 mitigates the catastrophic forgetting often observed when smaller models are finetuned on synthetic data from larger ones. As a result, it retains strong general language communication skills and explainable reasoning. o leverage these language skills for multi-agent traffic coordination, we introduce an asynchronous communication network that enables LLM agents to coordinate across multiple intersections via message-passing, mimicking the collaboration of human traffic agents.
In summary, our contributions are:
@@
  • Foundation model for TSC systems: We present the first LLM-based, general-purpose controller that can operate at any interaction without additional training, handling routine signaling, incident management, and emergency-vehicle prioritization with human-level reasoning.
  • Lightweight yet high-performing: Our two-stage RL training yields a 3B-parameter LLM that matches or exceeds much larger models (e.g., GPT-4o) and strong expert baselines, while remaining suitable for inference on edge devices.
@@
  • Human-like reasoning and communication: We achieve human-like reasoning for explainable TSC through self-iteration of the reinforced LLM. This reasoning, combined with language capabilities, is further utilized in our asynchronous communication network for effective coordination across multiple intersections.
@@
  • Extensive validation and field deployment: Evaluated on standard TSC benchmarks and out-of-distribution incident tasks, achieving stable state-of-the-art performance. In live deployment serving over 55,000 drivers daily, trials show a >5% reduction in average queues and a >50% reduction in operator workload for phase planning and incident response.
2 Related Works
@@
Traffic Signal Control. TSC is essential for traffic management. The FixedTime method, one of the earliest and most widely used approaches, relies on predetermined cycle lengths and phase allocations set by human experts for each intersection Yau et al. (2017); Serafini and Ukovich (1989). It is simple and stable but inefficient under dynamic traffic and requires substantial human effort Serafini and Ukovich (1989); Thunig et al. (2019). Maxpressure introduces adaptive control by prioritizing movements based on pressure, defined from queue lengths per direction Varaiya (2013); Mercader et al. (2020), yet it and FixedTime remain constrained by fixed rules that struggle with varying scenarios Rasheed et al. (2020); Wei et al. (2021); Abdulhai et al. (2003). Machine learning has produced RL-based TSC methods such as CoLight Wei et al. (2019), CosLight Ruan et al. (2024), and MPLight Chen et al. (2020), which improve performance in simulation but often fall short for real-world deployment Chen et al. (2022); Wei et al. (2021); Qadri et al. (2020). Recent work explores integrating large language models into TSC, but significant challenges remain before these approaches are practical for field use Lai et al. (2023); Yuan et al. (2025); Wang et al. (2025a, 2024); Wen et al. (2023); Zhao et al. (2025); Feng et al. (2025).
RL finetuning for LLM. LLMs often require fine-tuning to adapt to specific tasks or human preferences Wu et al. (2025); Minaee et al. (2024); Ding et al. (2025); Li et al. (2025). Instruction finetuning Chung et al. (2024); Dettmers et al. (2023) is widely used because it is simple and effective for task- and format-specific adaptation. RL finetuning, particularly RL from Human Feedback Ouyang et al. (2022), has seen less adoption for downstream use due to its complex pipeline and high compute costGao et al. (2023); Bai et al. (2022); Ramamurthy et al. (2022). Recent work shows LLMs can develop reasoning through interaction with RL environments, and Liu et al. (2024) reduces compute needs by using group policy-based optimization. RL finetuning offers practical advantages: (1) self-exploration reduces dependence on large labeled datasets, (2) it guides models to acquire capabilities rather than imitate or memorize outputs, and (3) KL-divergence constraints on updates help mitigate catastrophic forgetting by anchoring the policy to the pretrained distributionCao et al. (2024); Xu et al. (2025); Ke et al. (2025).
3 Preliminaries
Definition 3.1.
@@ 3 Preliminaries
4 Methodology
Figure 2: Introduction of the two-stage RL framework
@@ 4 Methodology
In this section, we describe the training pipeline for Traffic-R1. As shown in Figure 2, we utilize a two-stage RL framework comprising offline human-informed RL and online open-world RL to finetune the LLM. For each stage, we define distinct training datasets, environments, policy optimization methods, and reward designs to support their functionality. Besides, we propose an asynchronous communication network for reinforced LLM to support efficient multi-intersection coordination.
4.1 Human-Informed RL
@@
Existing LLMs for TSC are typically finetuned to imitate traditional RL models. This approach has two major flaws: first, finetuning LLMs to imitate traditional RL models confines their performance and generalization to the capabilities of the teacher RL model. Second, action trajectories produced by RL models through iterative optimization in simulated environments may not provide clear reasoning and logic for LLMs, as some RL decision policies and actions can be suboptimal or impractical for general TSC, focusing instead on narrow performance metrics. To address these limitations, we propose a human-informed RL finetuning stage that incorporates an expert-collaborative TSC dataset and offline policy optimization. This approach replaces the RL model teacher with real human traffic experts to guide the finetuning of the LLM.
4.1.1 Expert-Collaborative TSC Dataset
Our RL finetuning approach reduces the training data requirement from hundreds of thousands of samples typically needed for instruction finetuning to just thousands. This enables the creation of a TSC dataset with actions provided by human experts for each traffic scenario. We developed a 3,000-sample dataset using a human-in-the-loop pipeline. For each traffic scenario, a base model suggests an action. This action is then validated for effectiveness using the SUMO simulator and reviewed by two human traffic control experts. If an action is rejected, experts provide the correct one. This process, involving 11 experts, resulted in our final dataset. Crucially, we only include the final, validated actions, excluding any reasoning steps. This design encourages the model to develop its own reasoning capabilities during RL finetuning, rather than imitating an external thought process.
Figure 3: Expert-Collaborative Dataset Construction
4.1.2 Offline Policy Optimization